# MaNGA DynPop – VII. A Unified Bulge-Disk-Halo Model for Explaining Diversity in Circular Velocity Curves of 6000 Spiral and Early-Type Galaxies

### v1

@@ 4 Conclusions
This work is supported by the National Science Foundation of China (Grant No. 11821303 to SM). KZ acknowledges the support from the Shuimu Tsinghua Scholar Program of Tsinghua University. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
@@ Appendix A Dependence of the BDH model on the assumptions
We adopt a spherical exponential disk assumption, $V_{\rm disk}(r)=\sqrt{GM_{\rm disk}(r)/r}$, in the standard BDH model (Figure 5), which might not be a good approximation for LTGs. Here we test if the adopted assumptions affect the results. In Figure 7, we adopt the infinitely thin disk (Freeman, 1970) assumption for all galaxies (including ETGs and LTGs). Under this alternative extreme assumption, the circular velocity at radius $r$ (in the plane of disk) contributed from the disk component is
@@
where $M_{\rm\infty,disk}$ is the total mass of disk, $R_{\rm d}=R_{\rm e,disk}/1.678$ is the disk scale length, $y=(r/R_{\rm d})$, and $I_{\rm n}$ and $K_{\rm n}$ are the modified Bessel functions (Freeman, 1970). We find the predicted normalized $V_{\rm circ}$ profiles under the two extreme assumptions (spherical exponential disk or thin exponential disk) are nearly identical. In Figure 8, we fix the Sersic index of the bulge component to be 2.83, which is similar to Hernquist (1990) profile (see Appendix A and figure A1 in Vitral & Mamon, 2020), and find that adopting a universal Hernquist bulge can roughly predict the trend of CVCs (rising, flat or declining) but can not quantitatively reproduce the shape of CVCs in the inner region.
Figure 7: Similar to Figure 5, but adopt an alternative extreme assumption that all galaxies have infinitely thin exponential disk (Freeman, 1970).
Figure 8: Similar to Figure 5, but use a fixed $n_{\rm Ser,bulge}=2.83$ that corresponds to a Hernquist profile in the inner region (Appendix A and figure A1 in Vitral & Mamon, 2020). The trends of CVC shape changing with $\sigma_{\rm e}$ still exist, but BDH model predicted curves tend to overestimate at the low $\sigma_{\rm e}$ end and underestimate at the high $\sigma_{\rm e}$ end.

### v2

@@ 4 Conclusions
We thank the anonymous referee for the very useful comments. This work is supported by the National Science Foundation of China (Grant No. 11821303 to S.M.). K.Z. acknowledges the support from the Shuimu Tsinghua Scholar Program of Tsinghua University. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
@@ Appendix A Dependence of the BDH model on the assumptions
We adopt a spherical exponential disk assumption, $V_{\rm disk}(r)=\sqrt{GM_{\rm disk}(r)/r}$, in the standard BDH model (Figure 5), which might not be a good approximation for LTGs. Here we test if the adopted assumptions affect the results. In Figure 7, we adopt the infinitely thin disk (Freeman, 1970) assumption for all galaxies (including ETGs and LTGs). Under this alternative extreme assumption, the circular velocity at radius $r$ (in the plane of the disk) contributed from the disk component is
@@
where $M_{\rm\infty,disk}$ is the total mass of the disk, $R_{\rm d}=R_{\rm e,disk}/1.678$ is the disk scale length, $y=(r/R_{\rm d})$, and $I_{\rm n}$ and $K_{\rm n}$ are the modified Bessel functions (Freeman, 1970). We find the predicted normalized $V_{\rm circ}$ profiles under the two extreme assumptions (spherical exponential disk or thin exponential disk) are nearly identical. In Figure 8, we fix the Sersic index of the bulge component to be 2.83, which is similar to the Hernquist (1990) profile (see Appendix A and figure A1 in Vitral & Mamon, 2020), and find that adopting a universal Hernquist bulge can roughly predict the trend of CVCs (rising, flat or declining) but can not quantitatively reproduce the shape of CVCs in the inner region.
Figure 7: Similar to Figure 5, but adopting an alternative extreme assumption that all galaxies have infinitely thin exponential disks (Freeman, 1970).
Figure 8: Similar to Figure 5, but using a fixed $n_{\rm Ser,bulge}=2.83$ that corresponds to a Hernquist profile in the inner region (Appendix A and figure A1 in Vitral & Mamon, 2020). The trends of the CVC shape changing with $\sigma_{\rm e}$ still exist, but the BDH model-predicted curves tend to overestimate at the low $\sigma_{\rm e}$ end and underestimate at the high $\sigma_{\rm e}$ end.